OpenStudy (anonymous):

y = 5x + 6 y = −x − 7

OpenStudy (kidrah69):

\[ −x − 7 = 5x + 6\]

OpenStudy (kidrah69):

solve.

OpenStudy (anonymous):

13/-5?

OpenStudy (kidrah69):

hmmm not sure about the -5 :/ its supposed to be 6 how many solutions are there ?:)

OpenStudy (anonymous):

how did you get 6?

OpenStudy (kidrah69):

5x+x=6x

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

here's what i did maybe you could tell what im doing wrong -x - 7 = 5x +6 + 7 +7 -x = 5x + 13 then i added 5x to -x -4x = 13

OpenStudy (kidrah69):

it would be 6x.... the 5 is positive...

OpenStudy (kidrah69):

we are adding X

OpenStudy (anonymous):

but x is negative

OpenStudy (greencat):

ah hah!

OpenStudy (greencat):

you added 5x you shoulda subtracted

OpenStudy (greencat):

anyways that way is inconvenient

OpenStudy (greencat):

add x to each side; that's easier.

OpenStudy (anonymous):

how should i do it?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so -13 / 6x?

OpenStudy (anonymous):

but how do i do a fraction on a graph?

OpenStudy (greencat):

it is not -13/6x the answer is x=-13/6

OpenStudy (greencat):

The graph refers to another way to solve.

OpenStudy (greencat):

Two ways to solve: 1. Equate the two equations and solve for x. 2. Draw the graphs of both equations and see where they hit.

OpenStudy (anonymous):

oh i need two lol thanks

OpenStudy (anonymous):

wait so how would i do that?

OpenStudy (greencat):

Graph both. It should be simple since they are linear equations.

OpenStudy (anonymous):

okay thanks

OpenStudy (greencat):

Sidenote: If you are wondering how to find y using the first method... Just plug the calculated x value into the equation of your choice to calculate y.

OpenStudy (anonymous):

oh okay thanks